Practice 17: Semiconductor Contamination Information about product failure based on chip manufacturing process contamination level is given below. Find the probability of fallure. Probability Level of Probability of Failure Contamination of Level | 01 High 0.2 0.005 Not High 08
Answer #1

P(Failure)

= [P(Failure | Level of Contamination = High) * P(Level of Contamination = High)] + [P(Failure | Level of Contamination = Not High) * P(Level of Contamination = Not High)]

= (0.1*0.2) + (0.005*0.8)

= 0.024 (ans)